login | register

Projectionists Meeting @ 2012-01-30 18:00:00

Monday 30/01/12 (Spring week 4) 18:10 SSL

1. Welcome and Apologies for Absence
Present : Tim G (chair), Rob M, Neil A, Max H, Tristan B, Fergus C (late), Mike L (late)

In Attendance : Georgie R, Tim D, Andy M, Enrik N

Apologies : Yolanda C, Ed M Omar A, Alex W, Kath B, Rob G

2. Technical Officer's Report
The Technical Officer reported that all equipment appears to be in working order except the intermittent arm spring in P1. A new one will need to be sourced.

With regard the on-going Maintenance contract issue - Rob M asked whether we require remote monitoring. The response was that it depended on the cost of the monitoring on-top-of the basic contract versus what is gained from the monitoring.

Fergus Arrived.

3. Training
Given that the last few digital shows have been solo projected, it was decided that we have now reached a point when digital shows can be opened to trainees. The large amount of 'free' time during the shows should be taken advantage of for training.

Enrik's Snaps Test will be this Sunday. Rob M, Fergus C, Neil A and Tim G said they could attend for at least part of it. To the mailing list it shall go to see if we can drum-up more Qproj interest.

Tim G once again asked if we are ready to add more trainees to the scheme. It was decided that perhaps one more trainee can be added, but considering the lack of availability of slots in the next few weeks some shows should be PDFed for them.

4. Projection Procedure Changes
Tim G wanted to remind everyone that the busy show shut-down procedure must be initiated immediately following the end of the final trailer (see instructions), otherwise the instructions will not work.

Another reminder - to use the microphones - the receiver (which is now in the old data proj projection-port) must be turned on.

After a recommendation from a projectionist that aperture plates should be left in the projectors after every screening, it was decided that people were less likely to be muppets (and forget to choose the wrong aperture) if we continue as we have done.

Mike Arrived.

It was decided after some discussion that the start-up timetable sheet (which since digital start-up people have been forgetting to update) should be kept, but moved closer to the digital projector.

5. Time and Date of Next Meeting
The next meeting will be on Monday the 5th February in SSL at 18:10 and could well be Tim G's last meeting as chair!

6. AOB
After investigating why many of the films in the dump have not been collected, Andy noticed that several of the film labels had not been reversed. A reminder to all - films will not be collected if the labels are not reversed.

Meeting Ended at 18:49

Back to list